OpenStudy (babyx3boo):

Use the distributive property to simplify the expression: 4/7(21z-7)=___

OpenStudy (anonymous):

cancel as you go

OpenStudy (anonymous):

\[\frac{4}{7}\times 21=4\times 3\] for example

OpenStudy (babyx3boo):

84z-4 ?

OpenStudy (babyx3boo):

@satellite73

OpenStudy (anonymous):

what are the common factors in (21z-7) ?

OpenStudy (babyx3boo):

7 o.o

OpenStudy (anonymous):

oh no, should be a 12 out front, not \(84\)

OpenStudy (babyx3boo):

why a 12? @satellite73

OpenStudy (anonymous):

yes, so if you factor out a seven hat is left?

OpenStudy (babyx3boo):

never mind i get @satellite73

OpenStudy (babyx3boo):

cancelation

OpenStudy (anonymous):

\[\frac{4}{7}\times 21z=4\times 3z=12z\]

OpenStudy (babyx3boo):

so 12z-4

OpenStudy (anonymous):

sort of like the last one we did where you cancelled as you went along

OpenStudy (anonymous):

yeah that